# Hints for a Geon from Causal Dynamic Triangulations

### v2

@@ I Introduction
Investigating these questions within quantum gravity is the aim of this work. While the questions raised apply to any version of quantum gravity, it may be answered differently in each. In this work we exploit (causal) dynamical triangulations (CDT) [5, 10] which are accessible to non-perturbative simulations, feature a “long-distance“ de Sitter-like dynamics [5, 10], appear to be connected to asymptotic safety [11] as another candidate for quantum gravity [12, 13, 14], and allow to access to both the curvature [15, 16] and correlation functions [17, 18]. In particular the latter two are necessary ingredients for analyzing curvature correlators for signs of particle-like aspects.
@@ II Technical setup
For the creation of the configurations we use the code of [21]. For technical details of CDT simulations and notations, we refer to [5]. For the exploratory simulations at hand, we employed a single simulation point at the bare parameters $\Delta=0.6$ and $\kappa_{0}=2.2$, which exhibits a de Sitter-like space time [10], and thus a (classical) positive cosmological constant. Moreover, the lattice spacing, i. e. the (spatial) length of the simplices used to triangulate the space-time, has been estimated to be about $\approx 2.1$ Planck lengths [22], which we will take here at face value. We performed simulations on $N_{t}\times N_{\text{simp}}=60\times 80$k, $80\times 160$k, and $80\times 320$k, with $N_{t}$ the temporal extent of the foliated space-time (i.e. the number of time slices, or spacing between timelike distances) and $N_{\text{simp}}$ the number of (4,1)-simplices, which connect the time slices. Since we did not find any statistical significant dependence on the volume, we will show in the following only results from the largest volume [20]. We denote quantities at fixed foliation a time-slice, and everything on or in between two neighboring foliations a fat time-slice. In total, we calculated for each case 5000 (2000 for $N_{\text{simp}}=320$k) configurations, following standard procedures for thermalization and decorrelation [23]. We choose spatially spherical boundary conditions at fixed foliation and periodic boundary condition in foliation direction. We note that the effective obtained geometry is found [22, 24] to be four-dimensional spherical, and thus (Euclidean) de Sitter.
@@
Distance measures between non-neighboring points themselves are now non-trivial. Following [5, 26, 6], we take as a space-like distance measure the geodesic distance $d$ between dual vertices on a fixed fat foliation slice. This is obtained by employing a Dijkstra algorithm, which is run on the triangulation restricted to a certain fat foliation slice. Since the determination using all possible pairs is, of course, prohibitively expensive, we have uniformly sampled 100 (50 for $N_{\text{simp}}=80$k) pairs per fixed fat foliation slice for a stochastic estimator [27]. This is similar to the discussion in [18], in particular their remarks in Appendix B, which we also took into account.
@@
where the sums run over all points of the dual triangulation, i.e. the centers of simplices. As $Q$ does no longer depend on $\epsilon$ explicitly but only through the summation, it is possible to identify $p$ and $p^{\prime}$ by setting $\epsilon=0$, thus yielding a quantity measuring the distance between pairwise points in a sphere with respect to each other [18]. Thus, $Q$ is a measure of such distances, and triangulations with different local structure will in general yield locally different values of $Q$. In practice, we fixed $\delta=6$ for this work. However, when varying $\delta$ in some range around this value, we did not see any significant change [20].
@@
where the $q_{i}$ are positive, real constants, $R$ is the standard, continuum, curvature scalar. Thus, $Q$ is in this limit a measure for curvature. Indeed, this was confirmed to be a good approximation of the functional dependency on $\delta$ in [28] for CDT simulations. We confirmed this in the present work.
@@ III Defining a geon correlation function
It is well established [10, 24] that a single configuration, and even the average, in CDT simulations at the given simulation point do not resemble a flat Minkowski manifold, but a(n Euclidean) de Sitter space. This implies that it is impossible to speak about locations across configurations. In addition, the number of time-slices is usually not very large compared to the number of time-slices where the de Sitter space is almost static. It is a priori not clear, how the global space-time characteristics would affect the propagation of a gravitational state. We therefore opt for another approach.
@@ IV Results
Figure 1: Global properties as a function of cosmological time. The extent in simplices is measured by the cube root of all simplices with a vertex on a given time slice. Error bars are smaller than the plot points.
The global properties are shown in figure 1. As is visible, after introduction of the cosmological time, there is very little fluctuations in terms of the spatial size. As noted previously [5, 10], the generated universes are of very similar properties, allowing to indeed identify our configuration-wise cosmological time to play a meaningful global role. We also observe the characteristic transition of the (Euclidean) de Sitter universe from a very small object to a very large one, which in Minkowski signature would be an exponential (inflationary) phase. Only our periodic boundary condition limits this growth.
We note that the average size of the QRCS changes, from a smaller universe in the small-extent time to a larger value. This change is not smooth, and the fluctuations peak briefly close to the onset of the expansion. Surprisingly, fluctuations are large not in the small extent time period, but rather in the large volume period.
Figure 2: The inverse correlation function for $O=1$ with fits of type $a+b\cosh(m(s-s_{0}))$.
@@
Figure 3: The normalized correlation function for $O=\Delta Q$ with fits of type $a+b\exp(-ms)$ and their ±$\pm 1\sigma$ error band.
@@
The shape of the correlators remains roughly the same, and gets only more compressed with increasing cosmological time, and thus decreasing spatial extent of the universe. However, as Fig. 4 shows, the mass parameter, determining the exponential decay, is relatively stable, except during the phase of rapid expansion, where it quickly rises. It remains, however, approximately the same for both operators. This suggest that during the expansion of the Universe in our CDT simulations, the properties of geons changes drastically, but are more or less constant at times of constant extension. This pattern is the same for all volumes, and we observe no statistically significant volume dependence. The masses, averaged over $120\leq\tau<$ 12, are found to be 0.18(1), 0.16(1), 0.17(3) for the $Q$ correlator and 0.14(1), 0.17(1), 0.15(3) for the $Q^{2}$ correlator, for $N_{\text{simp}}$ being 80k, 160k, and 320k, respectively. It would correspond to a mass of $\sim 0.09M_{\text{Planck}}$.

### v3

@@ I Introduction
Investigating these questions within quantum gravity is the aim of this work. While the questions raised apply to any version of quantum gravity, it may be answered differently in each. In this work we exploit (causal) dynamical triangulations (CDT) [5, 10] which are accessible to non-perturbative simulations, feature a “long-distance“ de Sitter-like dynamics [5, 10], appear to be connected to asymptotic safety [11] as another candidate for quantum gravity [12, 13, 14], and allow to access both the curvature [15, 16] and correlation functions [17, 18]. In particular the latter two are necessary ingredients for analyzing curvature correlators for signs of particle-like aspects.
@@ II Technical setup
For the creation of the configurations we use the code of [21]. For technical details of CDT simulations and notations, we refer to [5]. For the exploratory simulations at hand, we employed a single simulation point at the bare parameters $\Delta=0.6$ and $\kappa_{0}=2.2$, which exhibits a de Sitter-like space time on average [10], and thus a (classical) positive cosmological constant. Moreover, the lattice spacing, i. e. the (spatial) length of the simplices used to triangulate the space-time, has been estimated to be about $\approx 2.1$ Planck lengths [22], which we will take here at face value. We performed simulations on $(N_{t},N_{\text{simp}})=(60,80$k)), (80,160(80,160k)), and (80,320(80,320k)), with $N_{t}$ the temporal extent of the foliated space-time (i.e. the number of time slices) and $N_{\text{simp}}$ the number of (4,1)-simplices, which connect the time slices. Since we did not find any statistical significant dependence on the volume, i. e. within $2-3\sigma$, we will show in the following only results from the largest volume [20]. We denote quantities at fixed foliation a time-slice, and everything on or in between two neighboring foliations a fat time-slice. In total, we calculated for each case 5000 (2000 for $N_{\text{simp}}=320$k) configurations, following standard procedures for thermalization and decorrelation [23]. We choose spatially spherical boundary conditions at fixed foliation and periodic boundary condition in foliation direction. We note that the effective obtained geometry is found [22, 24] to be four-dimensional spherical, and thus (Euclidean) de Sitter.
@@
[…]
@@
where the sums run over all points of the dual triangulation, i.e. the centers of simplices. As $Q$ does no longer depend on $\epsilon$ explicitly but only through the summation, it is possible to identify $p$ and $p^{\prime}$ by setting $\epsilon=0$, thus yielding a quantity measuring the distance between pairwise points in a sphere with respect to each other [18]. Thus, $Q$ is a measure of such distances, and triangulations with different local structure will in general yield locally different values of $Q$. In practice, we fixed $\delta=6$ for this work. A detailed analysis of the $\delta$-dependence will be given elsewhere [20]. However, different values of $\delta$ should be considered to define different operators, akin to differntely smeared operators in flat space-time [27]. Thus, long-distance physics should not be affected by the choice of $\delta$. We will give support that this is the case for our primary observable below. Other than that, the choice is here primarily technically motivated, as smaller $\delta$ have fewer points to average over the sphere, while larger ones become prohibitevly expensive in terms of computer time.
@@
where the $q_{i}$ are positive, real constants, and $R$ is the standard, continuum, curvature scalar. Thus, $Q$ is in this limit a measure for curvature. Indeed, this was confirmed to be a good approximation of the functional dependency on $\delta$ in [28] for CDT simulations. We confirmed this in the present work. Below, we will give also a result for correlators build from $R$, which was obtained by fitting for each measurement of $Q$ to (1) to determine measurmenets of $R$.
@@ III Defining a geon correlation function
It is well established [10, 24] that a single configuration, and even the average, in CDT simulations at the given simulation point do not resemble a flat Minkowski manifold, but are frequently close to a(n Euclidean) de Sitter space. This implies that it is impossible to speak about locations across configurations. In addition, the number of time-slices is usually not very large compared to the number of time-slices where the de Sitter space is almost static. It is a priori not clear, how the global space-time characteristics would affect the propagation of a gravitational state. We therefore opt for another approach.
@@ IV Results
Figure 1: Global properties as a function of cosmological time. The extent in simplices is measured by the cube root of all simplices with a vertex on a given fat time slice. Error bars are smaller than the plot points.
The global properties are shown in figure 1. As is visible there is very little fluctuations in terms of the spatial size, which is here given in terms of the cube root of simplicies per fat time slice [^1] If the universe would be spatially perfectly spherical, the maximum geodesic distance is around half the sphere, and about 3 times larger. In the figures (2) and (3) this is visible, as the maximal spatial geodesic distance plotted there is around twice as large.. Using [22], this implies a diameter of about 70 Planck lengths as maximum physical extent, and about 10 Planck lengths as minimal physical extent.
As noted previously [5, 10], the generated universes are of very similar properties, allowing to indeed identify our configuration-wise cosmological time to play a meaningful global role. We also observe the characteristic transition of the (Euclidean) de Sitter universe from a very small object to a very large one, which in Minkowski signature would be an exponential (inflationary) phase. We note that the average size of the QRCS changes, from a smaller universe in the small-extent time to a larger value. This change is not smooth, and the fluctuations peak briefly close to the onset of the expansion. Surprisingly, fluctuations are large not in the small extent time period, but rather in the large volume period.
Figure 2: The inverse correlation function for $O=1$ with fits of type $a+b\cosh(m(s-s_{0}))$. It is shown for three cosmological times, which can be compared to figure 1: $\tau=0$ (blue) is the largest extent, $\tau=22$ (orange) is the inflection point of the $Q$ change, and $\tau=32$ (red) is the point where the system settles on the smallest extent. The perpendicular lines give the size at the corresponding $\tau$ in terms of the cube root of simplices, for illustration of the relative sizes.
@@
Figure 3: The normalized correlation function for $O=\Delta Q$ with fits of type $a+b\exp(-ms)$ and their $\pm 1\sigma$ error band (top panel) at $\tau=0$. The same for the $O=R$, extracted using (1) for each measurement of pairs (bottom panel). The fit fixes the same mass as in the top panel, and only determines $a$ and $b$. We note that the first two points in distance could be affected by discretization errors, and therefore deviate from an exponential.
@@
The shape of the correlators remains roughly the same, and gets only more compressed with increasing cosmological time, and thus decreasing spatial extent of the universe. However, as Fig. 4 shows, the mass parameter, determining the exponential decay, is relatively stable, except during the phase of rapid expansion, where it quickly rises. It remains, however, approximately the same for both operators. This suggest that during the expansion of the Universe in our CDT simulations, the properties of geons changes drastically, but are more or less constant at times of (roughly) constant extension. This pattern is the same for all volumes, and we observe no statistically significant volume dependence. The masses, averaged over $0\leq\tau<12$, are found to be 0.18(1), 0.16(1), 0.17(3) for the $Q$ correlator and 0.14(1), 0.17(1), 0.15(3) for the $Q^{2}$ correlator, for $N_{\text{simp}}$ being 80k, 160k, and 320k, respectively. It would correspond to a mass of $\sim 0.09M_{\text{Planck}}$ using the result of [22].
We note that our system is relatively coarse, and thus the conversion to physical units may well have associated discretization artifacts. Only investigations along lines-of-constant physics can eventually clarify such a potential dependence. However, the fact that the result is within a few $\sigma$ independent on the maximal spatial extent suggests that the mass will eventually not be zero. Thus, the number should be considered for illustrative purpsoes only at the present time.